# Multi-virtual braid groups

### v1

@@ Abstract.
L. Kauffman (2024) introduced multi-virtual and symmetric multi-virtual braid groups, which are generalizations of the virtual braid group. We introduce multi-virtual pure and multi-virtual semi-pure braid groups, which are normal subgroups of index $n!$. We give a set of generators and defining relations for these groups, show that multi-virtual (symmetric multi-virtual) braid group is a semi-direct products of multi-virtual pure (symmetric multi-virtual pure) braid group and symmetric group. Also, we introduce multi-welded and multi-unrestricted braid groups and examines structure of three-strand 2-virtual braid group and some its subgroups and quotients. The paper concludes by outlining open problems and suggesting avenues for future research in this area.
@@ 1. Introduction
In this paper we are studying the following questions. Find generators and defining relations for pure and semi-pure multi-virtual braid groups. Define some quotients of these groups and find their generators and defining relations.
The paper is organized as follows. In Section 2, we recall some known definitions and known facts from braid theory. In Subsections 2.1 and 2.2 one can finds definitions and some known facts from classical and virtual braid theories. In particular, we give definitions and presentations for virtual pure and virtual semi-pure braid groups. In Subsection 2.3, we recall multi-virtual knot theory in brief.
In Section 3, we recall the Kauffman’s definitions of multi-virtual braid group and symmetric multi-virtual braid group. Further, we introduce multi-virtual pure braid group and multi-virtual semi-pure braid group. We give sets of generators and defining relations for these groups. Similar results we give for the symmetric multi-virtual braid group. In Section 4, we introduce multi-welded braid group and multi-unrestricted braid groups, we find their set of generators and defining relation. Also, we defined pure and semi-pure subgroups in these groups. In Section 5, we find presentations for 2-virtual braid group with 3-strand subgroups and presentations for some subgroups and quotients.
@@ 2.2. Virtual braid group
[…]
@@
[…]
@@
that is a group with one defining relation. This group is the flat virtual pure braid group and is the quotient of $VP_{3}$ by relations $\lambda_{ij}=\lambda_{ji}^{-1}$.
@@
[…]
@@ 2.3. Knot theory and multi-virtual knot theory
The famous Reidemeister’s theorem [22] says that the set of local moves $(r1)-(r3)$ of link’s diagrams, shown in the first row in Figure 2 generate isotopy of knots and links. Reidemeister showed that if we have two knots (links) in 3-dimensional space, then they are ambient isotopic if and only if corresponding diagrams for them can be obtained, one from the other, by a finite sequence of moves $(r1)-(r3)$.
@@
Two virtual link diagrams are virtually equivalent if one diagram can be transformed into the other by a sequence of classical and virtual Reidemeister moves as shown in Figure  2. The virtual Reidemeister moves are equivalent to a single move, the detour move, which is executed by selecting a segment of a component of the link diagram that contains no classical crossings (see Figure 3). After removing this segment, we may insert a new segment with no triple points and any double points result in a new virtual crossing. A virtual link is an equivalence class of virtual diagrams modulo the Reidemeister moves referring to classical crossings and the detour moves.
Figure 3. Detour move for a virtual crossings of type 00
L. Kauffman [14] introduced multi-virtual knot theory with many types of virtual crossings. A diagram of multi-virtual link is analogous to a diagram of a virtual link, but in this case we have many types of virtual crossings. A multi-virtual link is an equivalence class of multi-virtual diagrams modulo the Reidemeister moves referring to classical crossings and the detour moves, in which we transform an arc, which contains virtual crossings only one type (see Figure 3, in which we denote this type of virtual crossings by 00). Remark that the detour moves with over types of virtual crossings are forbidden (see Figure 4, where $\alpha\not=0$).
@@ 3. Multi-virtual braid group
In this section, we recall (see [14]) definitions of the multi-virtual and the symmetric multi-virtual braid groups. The motivation for introduction of these groups is the multi-virtual knot theory (see Subsection 2.3), which is a generalization of the classical knot theory.
Let $k\geq 1$, $k$-multiple virtual braid group on $n$ strands, $M_{k}VB_{n}$ is a group which is generated by elements
@@ III. Braid relations
$F1\colon\sigma_{i}\sigma_{i+1}\,\rho_{i}^{(\alpha)}=\rho_{i+1}^{(\alpha)}\,\sigma_{i}\sigma_{i+1}$, $10\leq\alpha\leq k-$ 1;
$F2\colon\sigma_{i+1}\sigma_{i}\,\rho_{i+1}^{(\alpha)}=\rho_{i}^{(\alpha)}\,\sigma_{i+1}\sigma_{i}$, $10\leq\alpha\leq k-$ 1;
@@
$\rho_{i}^{(\gamma)}\,\rho_{i+1}^{(\gamma)}\,\rho_{i}^{(\beta)}=\rho_{i+1}^{(\beta)}\,\rho_{i}^{(\gamma)}\,\rho_{i+1}^{(\gamma)}$,   $10<\gamma<\beta\leq k-$ 1.
@@ 3.1. Multi-virtual pure braid group
Define the map ¯$\bar{}\colon M_{k}VB_{n}\to\Lambda_{n}$ which takes an element $w\in M_{k}VB_{n}$ to its representative $\overline{w}$ from $\Lambda_{n}$. In this case the element $w\overline{w}^{-1}$ belongs to $M_{k}VP_{n}$. By Theorem 2.7 of [18] the group $M_{k}VP_{n}$ is generated by
@@
If $\alpha=0$, then all $s_{\lambda,\rho^{(0)}_{i}}=e$ and $s_{\lambda,\sigma_{i}}=\lambda(s_{e,\sigma_{i}})\lambda^{-1}=\lambda(\sigma_{i}\rho_{i})\lambda^{-1}=\lambda(\lambda_{i,i+1}^{-1})\lambda^{-1}$, which is equal to some $\lambda_{kl}$, by Lemma 2.1. These calculations are done in Theorem 1 [3].
@@
Since $s_{e,\rho^{(\alpha)}_{i}}=\rho^{(\alpha)}_{i}\rho^{(0)}_{i}$, $s_{\lambda,\rho^{(\alpha)}_{i}}=\lambda(\rho^{(\alpha)}_{i})\lambda^{-1}$, which is equal to some $\lambda^{(\alpha)}_{kl}$ by Lemma 3.4. Therefore, generators of the group $M_{k}VP_{n}$ are
@@
The remaining $r_{2,\lambda}$, $\lambda\in\Lambda_{n}$, can be obtained from this relation using conjugation by $\lambda^{-1}$ and it gives the same relation, by Lemma 2.1. We have obtained $\lambda_{ij}^{(0)}\lambda_{kl}^{(0)}=\lambda_{kl}^{(0)}\lambda_{ij}^{(0)}$, where indices are different.
@@
The remaining $r_{3,\lambda}$, $\lambda\in\Lambda_{n}$. We have obtained $\lambda_{ij}^{(\alpha)}\lambda_{kl}^{(\alpha)}=\lambda_{kl}^{(\alpha)}\lambda_{ij}^{(\alpha)}$, where indices are different. The other commutative relations give rises to the relations: $\lambda_{ij}^{(\alpha)}\lambda_{kl}^{(\beta)}=\lambda_{kl}^{(\beta)}\lambda_{ij}^{(\alpha)}$, $10\leq\alpha\leq\beta\leq k-$ 1. Combining all these relations, we obtain the relation (7).
@@
The remaining relations $r_{4,\lambda},\lambda\in\Lambda_{n}$, can be obtained from this relation using conjugation by $\lambda^{-1}$. By Lemma 2.1, we obtain (11).
@@
Conjugating this relation by all representatives from $\Lambda_{n}$, by Lemma 3.4, we obtain (12). Rest of the mixed relations of $M_{k}VB_{n}$ give rise of trivial relations. Therefore, the group $M_{k}VP_{n}$ is defined by the relations (7)-(12). ∎
@@ 3.2. Multi virtual semi–pure braid group
It is similar to the proof of Theorem 3.5. ∎
@@
We have decomposition $M_{k}VB_{n}=M_{k}VH_{n}\rtimes S_{n}$ and $S_{n}$ acts on $M_{k}VH_{n}$ by the rules
Lemma 3.9.
$a^{-1}x_{ij}^{(\alpha)}a=x_{(i)\bar{a},(j)\bar{a}}^{(\alpha)},$
@@ 3.3. Symmetric multi-virtual braid group
Let $k\geq 2$. Symmetric $k$-multiple virtual braid group $\widetilde{M_{k}VB}_{n}$ is a group which is quotient of $M_{k}VB_{n}$ by the forbidden relations $F3$. But in $\widetilde{M_{k}VB}_{n}$ the following relations are also forbidden:
$F1\colon\sigma_{i}\sigma_{i+1}\,\rho_{i}^{(\alpha)}=\rho_{i+1}^{(\alpha)}\,\sigma_{i}\sigma_{i+1}$, $10\leq\alpha\leq k-$ 1;
$F2\colon\sigma_{i+1}\sigma_{i}\,\rho_{i+1}^{(\alpha)}=\rho_{i}^{(\alpha)}\,\sigma_{i+1}\sigma_{i}$, $10\leq\alpha\leq k-$ 1;

### v2

@@ Abstract.
L. Kauffman (2024) introduced multi-virtual and symmetric multi-virtual braid groups, which are generalizations of the virtual braid group. We introduce multi-virtual pure and multi-virtual semi-pure braid groups, which are normal subgroups of index $n!$. We give a set of generators and defining relations for these groups, show that multi-virtual (symmetric multi-virtual) braid group is a semi-direct products of multi-virtual pure (symmetric multi-virtual pure) braid group and symmetric group. Also, we introduce multi-welded and multi-unrestricted braid groups and some its subgroups and quotients. The paper concludes by outlining open problems and suggesting avenues for future research in this area.
@@ 1. Introduction
In this paper we provide generators and defining relations for pure and semi-pure multi-virtual braid groups and we define some new quotients of these groups, showing generators and defining relations.
The paper is organized as follows. In Section 2, we recall some known definitions and facts in (virtual) braid theory. In particular, we give definitions and presentations for virtual pure and virtual semi-pure braid groups. In Subsection 2.3, we recall multi-virtual knot theory in brief.
In Section 3, we recall the Kauffman’s definitions of multi-virtual braid group and symmetric multi-virtual braid group. Further, we introduce multi-virtual pure braid group and multi-virtual semi-pure braid group. We give sets of generators and defining relations for these groups. Similar results we give for the symmetric multi-virtual braid group. In Section 4, we introduce multi-welded braid group and multi-unrestricted braid groups, we find their set of generators and defining relation. Also, we defined pure and semi-pure subgroups in these groups.
@@ 2.2. Virtual braid group
The virtual pure braid group $VP_{n}$, $n\geq 2$, was introduced in [3] as the kernel of the homomorphism $VB_{n}\to S_{n}$, $\sigma_{i}\mapsto\rho_{i}$, $\rho_{i}\mapsto\rho_{i}$ for all $i=1,2,\ldots,i-1$. It was proved that $VP_{n}$ admits a presentation with the generators $\lambda_{ij},\ 1\leq i\neq j\leq n,$ and the following relations:
@@
The flat virtual braid group $FVB_{n}$ is the quotient of $VP_{n}$ by the relations $\sigma_{i}^{2}=e$ for all $i$. It contains the group of flat virtual pure braids $FVP_{n}$ that is generated by elements $\lambda_{ij},\ 1\leq i<j\leq n,$ and the map $\lambda_{ji}\mapsto\lambda_{ij}^{-1}$, $1\leq i<j\leq n,$ defines a homomorphism of $VP_{n}$ onto $FVP_{n}$. See Figure 3
Figure 2. Geometric interpretations of generators $\lambda_{ij}$ and $\lambda_{ji}$
@@
with one defining relation, is a subgroup of group $VP_{3}$. Since, there is a natural retraction $f:VP_{3}\to FVP_{3}$ such that $f\!\mid_{FVP_{3}}$ is the identity map.
This group is the flat virtual pure braid group and is the quotient of $VP_{3}$ by relations $\lambda_{ij}=\lambda_{ji}^{-1}$.
@@
where distinct letters stand for distinct indices. This presentation was found in [21] (see also [4]). We have decomposition $VB_{n}=VH_{n}\rtimes S_{n}$, where $S_{n}=\langle\rho_{1},\dots,\rho_{n-1}\rangle$ acts on the set $\{x_{i,j}\,,\;1\leq i\not=j\leq n\}$ by permutation of indices and we have the following analogous of Lemma 2.1.
@@ 2.3. Knot theory and multi-virtual knot theory
The famous Reidemeister’s theorem [22] says that the set of local moves (r1)(r3)(r1)-(r3) of link’s diagrams, shown in the first row in Figure 3 define equivalent relation on link’s diagrams. Reidemeister showed that if we have two knots (links) in 3-dimensional space, then they are ambient isotopic if and only if corresponding diagrams for them can be obtained, one from the other, by a finite sequence of moves (r1)(r3)(r1)-(r3).
@@
Two virtual link diagrams are virtually equivalent if one diagram can be transformed into the other by a sequence of classical and virtual Reidemeister moves as shown in Figure  3. The virtual Reidemeister moves are equivalent to a single move, the detour move, which is executed by selecting a segment of a component of the link diagram that contains no classical crossings (see Figure 4). After removing this segment, we may insert a new segment with no triple points and any double points result in a new virtual crossing. A virtual link is an equivalence class of virtual diagrams modulo the Reidemeister moves referring to classical crossings and the detour moves.
Figure 4. Detour move for a virtual crossings of type 0
Multi-virtual knot theory is a generalization of virtual knot theory in which a diagram may contain several distinct types of virtual crossings instead of a single kind. A multi-virtual knot diagram is an immersion of a circle in the plane whose double points are either classical crossings or virtual crossings labeled by a symbol $\alpha$, where $\alpha$ is simply a label taken from a fixed finite set (one may take it as an index set, e.g., integers, if desired). Equivalence is generated by the usual classical Reidemeister moves together with virtual Reidemeister–type moves that preserve the labels of virtual crossings. In particular, detour moves are allowed only when they respect the types (labels) of the virtual crossings involved. For more clarity see Figure 4 and 5. This framework reduces to ordinary virtual knot theory when all virtual crossings share one type. The motivation for this generalization is that multiple virtual types encode additional combinatorial and topological data, which allows finer distinctions between knots and supports the construction of stronger invariants. Kauffman developed this theory in the course of addressing the Penrose evaluation problem for colorings of arbitrary trivalent graphs.
@@ 3. Multi-virtual braid group
In this section, we recall (see [14]) definitions of the multi-virtual and the symmetric multi-virtual braid groups. The introduction of these groups is evidently motivated by multi-virtual knot theory recalled above.
Let $\geq 1, kk$-multiple virtual braid group on $n$ strands, $M_{k}VB_{n}$ is a group which is generated by elements
@@ III. Braid relations
$F1\colon\sigma_{i}\sigma_{i+1}\,\rho_{i}^{(\alpha)}=\rho_{i+1}^{(\alpha)}\,\sigma_{i}\sigma_{i+1}$, $0\leq\alpha\leq k-1$;
$F2\colon\sigma_{i+1}\sigma_{i}\,\rho_{i+1}^{(\alpha)}=\rho_{i}^{(\alpha)}\,\sigma_{i+1}\sigma_{i}$, $0\leq\alpha\leq k-1$;
@@
$\rho_{i}^{(\gamma)}\,\rho_{i+1}^{(\gamma)}\,\rho_{i}^{(\beta)}=\rho_{i+1}^{(\beta)}\,\rho_{i}^{(\gamma)}\,\rho_{i+1}^{(\gamma)}$,   $0<\gamma<\beta\leq k-1$.
@@ 3.1. Multi-virtual pure braid group
Define the map $\bar{}\colon M_{k}VB_{n}\to\Lambda_{n}$ which takes an element $w\in M_{k}VB_{n}$ to its representative $\overline{w}$ from $\Lambda_{n}$. In this case the element $w\overline{w}^{-1}$ belongs to $M_{k}VP_{n}$. By Theorem 2.7 of [18] the group $M_{k}VP_{n}$ is generated by
@@
If $\alpha=0$, then all $s_{\lambda,\rho^{(0)}_{i}}=e$ and $s_{\lambda,\sigma_{i}}=\lambda(s_{e,\sigma_{i}})\lambda^{-1}=\lambda(\sigma_{i}\rho_{i})\lambda^{-1}=\lambda(\lambda_{i,i+1}^{-1})\lambda^{-1}$, which is equal to some $\lambda^{(0)}_{kl}$, by Lemma 2.1. These calculations are done in Theorem 1 [3].
@@
Since $s_{e,\rho^{(\alpha)}_{i}}=\rho^{(\alpha)}_{i}\rho^{(0)}_{i}$, $s_{\lambda,\rho^{(\alpha)}_{i}}=\lambda(\rho^{(\alpha)}_{i}\rho^{(0)}_{i})\lambda^{-1}$, which is equal to some $\lambda^{(\alpha)}_{kl}$ by Lemma 3.4. Therefore, generators of the group $M_{k}VP_{n}$ are
@@
The remaining $r_{2,\lambda}$, for all $\lambda\in\Lambda_{n}$, can be obtained from this relation using conjugation by $\lambda^{-1}$ and it gives the same relation, by Lemma 2.1. We have obtained $\lambda_{ij}^{(0)}\lambda_{kl}^{(0)}=\lambda_{kl}^{(0)}\lambda_{ij}^{(0)}$, where indices are different.
@@
The remaining $r_{3,\lambda}$, for all $\lambda\in\Lambda_{n}$.
$r_{3,\lambda}=\tau(r_{3})$ $=s_{\lambda,\rho_{i}^{(\alpha)}}s_{\lambda\rho_{i},\rho_{j}^{(\alpha)}}s_{\lambda\rho_{i}\rho_{j},\rho_{i}^{(\alpha)}}s_{\lambda\rho_{i}\rho_{j}\rho_{i},\rho_{j}^{(\alpha)}}$
$=(\lambda(\lambda_{i,i+1}^{(\alpha)})^{-1}\lambda^{-1})(\lambda((\lambda_{j,j+1}^{(\alpha)})^{-1}\lambda^{-1})(\lambda((\lambda_{i,i+1}^{(\alpha)})^{-1}\lambda^{-1})(\lambda((\lambda_{j,j+1}^{(\alpha)})^{-1}\lambda^{-1})$
$=\lambda(\lambda_{i,i+1}^{(\alpha)})^{-1}(\lambda_{j,j+1}^{(\alpha)})^{-1}(\lambda_{i,i+1}^{(\alpha)})^{-1}(\lambda_{j,j+1}^{(\alpha)})^{-1}\lambda^{-1}.$
By Lemma 3.4, we have obtained $\lambda_{ij}^{(\alpha)}\lambda_{kl}^{(\alpha)}=\lambda_{kl}^{(\alpha)}\lambda_{ij}^{(\alpha)}$, where indices are different. The other commutative relations give rises to the relations: $\lambda_{ij}^{(\alpha)}\lambda_{kl}^{(\beta)}=\lambda_{kl}^{(\beta)}\lambda_{ij}^{(\alpha)}$, $0\leq\alpha\leq\beta\leq k-1$. Combining all these relations, we obtain the relation (7).
@@
The remaining relations $r_{4,\lambda}$, for all $\lambda\in\Lambda_{n}$, can be obtained from previous relation using conjugation by $\lambda^{-1}$. By Lemma 2.1, we obtain (11).
@@
Conjugating this relation by all representatives from $\Lambda_{n}$, by Lemma 3.4, we obtain (12). Rest of the mixed relations of $M_{k}VB_{n}$ become trivial relations. Therefore, the group $M_{k}VP_{n}$ is defined by the relations (7)-(12). ∎
@@
Example 3.7.
To illustrate definitions and results of this sections, we consider 2-virtual 3-strand braid group. As a rule, 3-strand braid groups are different from the braid groups with larger number of strands and are therefore studied separately (see [6], [8], [12], [16]).
By the definition, the group $MVB_{3}=M_{2}VB_{3}$ and it is generated by 6 elements:
$\sigma_{1},~~\sigma_{2},~~\rho_{1},~~\rho_{2},~~\tau_{1},~~\tau_{2},$
and is defined by relations
$\rho_{1}^{2}=\rho_{2}^{2}=\tau_{1}^{2}=\tau_{2}^{2}=e,$
The kernel $\ker\varphi_{3}$ of the homomorphism $\varphi_{3}\colon MVB_{3}\to S_{3}$, which is defined on the generators:
$\sigma_{i}\mapsto\rho_{i},~~\tau_{i}\mapsto\rho_{i},~~\rho_{i}\mapsto\rho_{i},~~i=1,2.$
The $\ker\varphi_{3}$ is 2-virtual pure braid group $MVP_{3}$. This group is generated by elements
$\mu_{12}=\rho_{1}\tau_{1},~~\mu_{23}=\rho_{2}\tau_{2},~~\mu_{13}=\rho_{2}(\rho_{1}\tau_{1})\rho_{2}=\rho_{2}\mu_{12}\rho_{2},$
and is defined by the relations
$\mu_{12}\mu_{13}\mu_{23}=\mu_{23}\mu_{13}\mu_{12}.$
@@ 3.2. Multi virtual semi–pure braid group
It is similar to the proof of Theorem 3.5. Define the map $\bar{}\colon M_{k}VB_{n}\to\Lambda_{n}$ which takes an element $w\in M_{k}VB_{n}$ to its representative $\overline{w}$ from $\Lambda_{n}$. In this case the element $w\overline{w}^{-1}$ belongs to $M_{k}VH_{n}$. By Theorem 2.7 of [18] the group $M_{k}VH_{n}$ is generated by
where $\alpha=0,1,\ldots,k-1$.
If $\alpha=0$, then all $s_{\lambda,\rho^{(0)}_{i}}=e$ and $s_{\lambda,\sigma_{i}}=\lambda(s_{e,\sigma_{i}})\lambda^{-1}=\lambda(\sigma_{i})\lambda^{-1}=\lambda(x^{(0)}_{i,i+1})\lambda^{-1}$, which is equal to some $x^{(0)}_{kl}$, by Lemma 2.3. These calculations are done in Proposition 17 [4].
Now, consider the generators such that $\alpha\neq 0$.
Since $s_{e,\rho^{(\alpha)}_{i}}=\rho^{(\alpha)}_{i}\rho^{(0)}_{i}$, $s_{\lambda,\rho^{(\alpha)}_{i}}=\lambda(\rho^{(\alpha)}_{i}\rho^{(0)}_{i})\lambda^{-1}$, which is equal to some $x^{(\alpha)}_{kl}$ by Lemma 3.8. Therefore, generators of the group $M_{k}VH_{n}$ are
$x^{(0)}_{kl},\text{ for }1\leq k\neq l\leq n\text{ and }x^{(\beta)}_{ij},~~1\leq i<j\leq n,~~1\leq\beta\leq k-1.$
To find the defining relations of $M_{k}VH_{n}$, we define a rewriting process $\tau$. It helps to rewrite a word $u$ to $\tau(u)$, where $u$ is written in the generators of $M_{k}VB_{n}$ but represents an element of $M_{k}VH_{n}$ and $\tau(u)$ is a word written in the generators of $M_{k}VH_{n}$. Let us associate to reduce word
the word
in the generators of $M_{k}VH_{n}$, where $k_{j}$ is the $(j-1)^{th}$ initial segment of the word $u$ if $\epsilon_{j}=1$, and a representative of the $j$-th initial segment of $u$ if $\epsilon_{j}=-1$.
By Theorem 2.9 in [18], the group $M_{k}VH_{n}$ is defined by the relations
where $r_{\mu}$ is a defining relation of $M_{k}VB_{n}$.
Let us consider relations of $M_{k}VB_{n}$ stated in Section 3.
Consider the $r=\sigma_{i}\sigma_{j}\sigma_{i}^{-1}\sigma_{j}^{-1}$, $|i-j|\geq 2$. Then
$\tau(r)$ $=s_{e,\sigma_{i}}s_{\rho_{i},\sigma_{j}}s^{-1}_{\rho_{i}\rho_{j}\rho_{i},\sigma_{i}}s^{-1}_{\rho_{i}\rho_{j}\rho_{i}\rho_{j},\sigma_{j}}$
$=(x^{(0)}_{i,i+1})^{-1}(x^{(0)}_{j,j+1})^{-1}(x^{(0)}_{i,i+1})(x^{(0)}_{j,j+1})$.
The remaining relations $\lambda\tau(r)\lambda^{-1}$, for all $\lambda\in\Lambda_{n}$, can be obtained from previous relation using conjugation by $\lambda^{-1}$ and it gives the same relation, by Lemma 2.3. We have obtained $x_{ij}^{(0)}x_{kl}^{(0)}=x_{kl}^{(0)}x_{ij}^{(0)}$, where indices are different.
We can work on other relations of $M_{k}VB_{n}$ similarly as done in Theorem 3.5. ∎
@@
We have decomposition $M_{k}VB_{n}=M_{k}VH_{n}\rtimes S_{n}$ and $S_{n}$ acts on $M_{k}VH_{n}$ by the rules.
@@ 3.3. Symmetric multi-virtual braid group
Let $k\geq 2$. Symmetric $k$-multiple virtual braid group $\widetilde{M_{k}VB}_{n}$ is a group which is quotient of $M_{k}VB_{n}$ by the forbidden relations $F3$.
But in $\widetilde{M_{k}VB}_{n}$ the following relations are also forbidden:
$F1\colon\sigma_{i}\sigma_{i+1}\,\rho_{i}^{(\alpha)}=\rho_{i+1}^{(\alpha)}\,\sigma_{i}\sigma_{i+1}$, $0\leq\alpha\leq k-1$;
$F2\colon\sigma_{i+1}\sigma_{i}\,\rho_{i+1}^{(\alpha)}=\rho_{i}^{(\alpha)}\,\sigma_{i+1}\sigma_{i}$, $0\leq\alpha\leq k-1$;